# HIF: The hypergraph interchange format for higher-order networks

### v1

@@ V.1.3 HyperNetX
For HyperNetX, we demonstrate the usefulness of analyzing the main component of the publications hypergraph using the $s$-closeness centrality measure for $s=1$, visualized as an Euler diagram in Fig. 9(c). We did not use the full hypergraph for this centrality analysis as it is too dense and disconnected to visually discern hyperedge centrality. The $s$-closeness centrality measure shows the closeness of each hyperedge (in this case, a publication) based on the shared authors. The color scale indicates the centrality value, with warmer colors (yellow) representing hyperedges that are more central, meaning they share more authors with other publications. This analysis can highlight influential publications that are central to the collaborative structure of this publication hypergraph.
V.1.4 SimpleHypergraphs.jl
@@
For the analysis with SimpleHypergraphs.jl we have selected the largest connected component of the citation hypergraph that has $108$ hypervertices and $30$ hyperedges. Next, we use library’s functionality hypergraph modularity-based clustering introduced in Ref. [60] to identify clusters within the citation hypergraph. Subsequently, we project the hypergraph to a pairwise projection where every hyperedge becomes a clique. We visualize the graph in Fig. 9(d), scaling node sizes by their degrees and using colors to denote community labels. All communities with a single node are colored in black.
V.1.5 XGI
@@
A strength of the XGI software package is its statistics interface, allowing practitioners to easily visualize, output as common Python data types, and compute statistics of node and edge properties. In Fig. 9(e), we see a hypergraph with its nodes colored and sized according to its degree and the average degree of its neighbors. XGI can also easily filter datasets by hyperedge size according to the framework in Ref. [61]. Another strength of the XGI software package is its ability to quickly remove data artifacts with its cleanup method which has the ability to remove singletons, isolated nodes, multi-edges, and disconnected components. In Fig. 9(e), we see the largest component of the hypergraph without any multi-edges, isolated nodes, or singletons.
VI Discussion
@@
As part of this effort, we surveyed the landscape of open-source software for existing standards and potential adopters, reaching out to more than 20 open-source projects with support for hypergraphs and simplicial complexes. Our exploration validated the hypothesis that the community of higher-order networks needs an open standard and that the HIF design supports many use cases. Future work will include integration with more software libraries to improve its reach in the higher-order network science community.
@@
Lastly, while the HIF standard directly supports undirected and directed hypergraphs and simplicial complexes, support for additional higher-order representations will make HIF more widely useful. Examples include hypergraphs with ordered and partially ordered sets (posets) which are useful in legal analysis [62] and the Science of Science. Several types of higher-order networks are implicitly supported by the HIF standard: signed hypergraphs which assign a sign +/- to each hyperedge; multilayer hypergraphs which assign nodes and hyperedges to different layers representing, for example, modes of transportation, relationship types, or social media platforms; and temporal hypergraphs where nodes and hyperedges have associated time intervals in which they are active. The specific structural information necessary for each of these representations can be stored in the attrs field, but the schema excludes these as expected JSON fields. In the future, adding explicit support for these fields will improve compatibility with libraries such as ASH [23], Raphtory [63], and Reticula [28]. In addition, we recommend that when releasing new higher-order datasets that an HIF-compliant file is included in the release (for example, see Ref. [64]).
@@
Data Availability
All datasets, resources, documentation, and schemas are available on GitHub footnotemark: . The stable schema is available on Zenodo [65] and subsequent stable versions will be released on Zenodo.
Contribution statement.
Author order is alphabetical. N.W.L. led the project. All authors designed the HIF standard and wrote the manuscript.
Funding.
N.W.L. acknowledges support from the University of Virginia Prominence-to-Preeminence (P2PE) STEM Targeted Initiatives Fund, SIF176A Contagion Science.
This work acknowledges support from Pacific Northwest National Laboratory with information release number PNNL-SA-213514

### v2

@@ 5.1.3 HyperNetX
For HyperNetX, we demonstrate the usefulness of analyzing the main component of the publications hypergraph using the $s$-closeness centrality measure for $s=1$, visualized as an Euler diagram in Figure 9(c). We did not use the full hypergraph for this centrality analysis as it is too dense and disconnected to visually discern hyperedge centrality. The $s$-closeness centrality measure shows the closeness of each hyperedge (in this case, a publication) based on the shared authors. The color scale indicates the centrality value, with warmer colors (yellow) representing hyperedges that are more central, meaning they share more authors with other publications. This analysis can highlight influential publications that are central to the collaborative structure of this publication hypergraph.
5.1.4 SimpleHypergraphs.jl
@@
For the analysis with SimpleHypergraphs.jl we have selected the largest connected component of the citation hypergraph that has $108$ vertices and $30$ hyperedges. Next, we use library’s functionality hypergraph modularity-based clustering introduced in kaminski_clustering_2019 to identify clusters within the citation hypergraph. Subsequently, we project the hypergraph to a pairwise representation where every hyperedge becomes a clique. We visualize the graph in Figure 9(d), scaling node sizes by their degrees and using colors to denote community labels. All communities with a single node are colored in black.
5.1.5 XGI
@@
A strength of the XGI software package is its statistics interface, allowing practitioners to easily visualize, output as common Python data types, and compute statistics of node and edge properties. In Figure 9(e), we see a hypergraph with its nodes colored and sized according to its degree and the average degree of its neighbors. XGI can also easily filter datasets by hyperedge size according to the framework in landry_filtering_2024. Another strength of the XGI software package is its ability to quickly remove data artifacts with its cleanup method which has the ability to remove singletons, isolated nodes, multi-edges, and disconnected components. In Figure 9(e), we see the largest component of the hypergraph without any multi-edges, isolated nodes, or singletons.
6 Discussion
@@
As part of this effort, we surveyed the landscape of open-source software for existing standards and potential adopters, reaching out to more than 20 open-source projects with support for hypergraphs and abstract simplicial complexes. Our exploration validated the hypothesis that the community of higher-order networks needs an open standard and that the HIF design supports many use cases. Future work will include integration with more software libraries to improve its reach in the higher-order network science community.
@@
Lastly, while the HIF standard directly supports undirected and directed hypergraphs and abstract simplicial complexes, support for additional higher-order representations will make HIF more widely useful. Examples include hypergraphs with ordered and partially ordered sets (posets) which are useful in legal analysis coupette_legal_2024 and the Science of Science. Several types of higher-order networks are implicitly supported by the HIF standard: signed hypergraphs which assign a sign +/- to each hyperedge; multilayer hypergraphs which assign nodes and hyperedges to different layers representing, for example, modes of transportation, relationship types, or social media platforms; and temporal hypergraphs where nodes and hyperedges have associated time intervals in which they are active. The specific structural information necessary for each of these representations can be stored in the attrs field, but the schema excludes these as expected JSON fields. In the future, adding explicit support for these fields will improve compatibility with libraries such as ASH failla_attributed_2023, Raphtory steer_raphtory_2024, and Reticula badie-modiri_reticula_2023. In addition, we recommend that when releasing new higher-order datasets that an HIF-compliant file is included in the release [for example, see smith_blue_2025a].
@@
Funding Statement
N.W.L. acknowledges support from the University of Virginia Prominence-to-Preeminence (P2PE) STEM Targeted Initiatives Fund, SIF176A Contagion Science. This work acknowledges support from Pacific Northwest National Laboratory with information release number PNNL-SA-217265. Pacific Northwest National Laboratory is a multiprogram national laboratory operated for the US Department of Energy (DOE) by Battelle Memorial Institute under Contract No. DE-AC05-76RL01830.
Competing Interests
None
Data Availability Statement
All datasets, resources, documentation, and schemas are available on GitHub footnotemark: . The stable schema is available on Zenodo zenodo_2025 and subsequent stable versions will be released on Zenodo.
Ethical Standards
The research meets all ethical guidelines, including adherence to the legal requirements of the study country.
Contribution statement
Author order is alphabetical.
Conceptualization: C.A.J., N.W.L., B.P., P.S.
Methodology: M.C., C.A.J., N.W.L., Q.F.L., A.M., J.P., B.P., P.S.
Project Administration and Supervision: N.W.L.
Software: M.C., N.W.L., Q.F.L., A.M., J.P., B.P., P.S.
Writing – Original Draft: M.C., C.A.J., N.W.L., Q.F.L., A.M., J.P., B.P., P.S.
Writing – Review & Editing: M.C., C.A.J., N.W.L., Q.F.L., A.M., J.P., B.P., P.S.
Visualization: M.C., N.W.L., Q.F.L., A.M., J.P., P.S.